OpenStudy (howard-wolowitz):

Find the sum of the geometric sum: 12+36+108+ ... + 78,732

OpenStudy (howard-wolowitz):

A. 354,288

OpenStudy (howard-wolowitz):

B. 118,092

OpenStudy (howard-wolowitz):

C. 39,360

OpenStudy (howard-wolowitz):

D. 1,062,876

OpenStudy (howard-wolowitz):

HELP

OpenStudy (howard-wolowitz):

@phi

OpenStudy (howard-wolowitz):

phi guess what

OpenStudy (mathmath333):

do u know the formula for sum of geometric series ?

OpenStudy (howard-wolowitz):

no i dont math man

OpenStudy (howard-wolowitz):

do you?

OpenStudy (mathmath333):

it is \(\large \begin{align} \color{black}{Sn=\dfrac{a(1-r^n)}{1-r}\\~\\ }\end{align}\) find \(a=\)first term \(r=\dfrac{2^{nd}term}{1^{st}term}\)=common ratio \(n\) =number of terms

OpenStudy (howard-wolowitz):

what did you get

OpenStudy (howard-wolowitz):

im guessiong option d

OpenStudy (mathmath333):

i dont know answer

OpenStudy (howard-wolowitz):

my experience tells me that it is that A option

OpenStudy (mathmath333):

u can find that by finding a,r and n.

OpenStudy (howard-wolowitz):

yes but i wana compare answers

OpenStudy (howard-wolowitz):

you get yous and I'll tell you mine!

OpenStudy (howard-wolowitz):

Find the sum of the geometric sum: 12+36+108+ ... + 78,732